an unbalanced force of 30N began to act on a resting body of 2 kg. What distance does the body travel during the first ten seconds of the movement?


1
Expert's answer
2021-04-29T10:43:06-0400
"s=0.5at^2\\s=0.5\frac{F}{m}t^2\\s=0.5\frac{30}{2}(10)^2=750\ m"